# Quantifying Uncertainty in Error Consistency: Towards Reliable Behavioral Comparison of Classifiers

### v1

@@ Appendix C Proof of Ceiling Effects
In Section 2, we claim that as soon as one classifier achieves perfect accuracy on the investigated sequence of trials, the EC of the pair will be 0 irrespective of the accuracy of the other one. Here, we proof this claim, which is easiest to derive from a confusion matrix as shown in Table 1. The observed agreement is found on the diagonal, $p_{obs}=a+d$, while the expected agreement is what would be on the diagonal if observers were independent ($p_{exp}=(a+b)(a+c)+(c+d)(b+d)$). For the proof, consider without loss of generality that classifier 2 was perfect, i.e. $a+c=1$.
@@
Appendix D Effect of EC on CI size
For Figure 3, we have chosen a ground-truth EC of 0.5 to demonstrate how the EC between two classifiers depends on their accuracy and the number of trials on which EC is measured. However, the choice of the ground-truth EC influences the CI size as well, which we demonstrate in Figure 6. We again generate data using our copy-model, and plot the ground-truth EC (input of the model) against the distribution of empirical ECs (calculated on model outputs). To better show the changes in CI size, we plot the delta between empirical and ground-truth EC rather than the absolute EC. Note how the CI bounds are not symmetric and largest at a ground-truth EC of 0.5.
@@ Appendix D Effect of EC on CI size
Appendix E Generating Synthetic Brain-Score Data
For the analysis conducted in Section 4.2, we estimate the size of confidence intervals surrounding models on the public Brain-Score benchmark. To estimate a model’s EC to humans within one condition $c$, we would technically have to create one set of binary trials that has the property of having the desired average EC to the $n$ human observers of $\kappa_{real,c}$. To simplify, we drop this dependence and generate $n$ independent sequences of trials, one for each human who saw images in this condition. Generating only one sequence of trials would drastically overestimate the variance, which is reduced by taking the mean. For each of these sequences, we need three inputs for the copy-model: The desired kappa (which is given by the Brain-Score data after removing the ceiling correction), the accuracy of the humans (which is given by the model-vs-human data) and the accuracy of the model in this condition. The latter is unknown, but bounded mathematically. We obtain the CI sizes in Figure 5 by selecting the accuracy in the middle of the bounds, which seems reasonable. Selecting the lower bounds of the accuracy leads to the most optimistic estimate of CI size, which we plot in Figure 7.
Appendix F Details on Model-VS-Human results
@@
Another idiosyncracy of the model-vs-human benchmark is that while the final scores (which are averages across the different corruptions) are stable, the uncertainty within each corruption can be quite large. In Figure 8, we plot the CIs surrounding the EC values obtained for the phase-scrambling corruption as an illustration of this point. In Figure 9 we plot the pairwise rank correlations between all 17 corruptions, revealing that some, but not all corruptions are highly correlated, i.e. imply similar model rankings.
@@
In Figure 2b, we explain how in many of the conditions within model-vs-human, the accuracy mismatch between models and human participants is so large that $k_{max}$ takes on very small values. One might wonder how model-vs-human scores would change if one excluded conditions not as prescribed by the authors, but based on this mismatch. However, this would render the selection of conditions dependent on the model. Hence, the benchmark would no longer present a fair comparison of models, and final scores would no longer be comparable.
Figure 7: Most optimistic Confidence Intervals for Brain-Score ECs. Like in Figure 5, we plot the 95% confidence interval of the EC of the top 30 Brain-Score models, but assume the lowest mathematically possible model accuracy. The true CI size for these models probably looks more like Figure 5, but this is a lower bound on their size.
@@
Figure 8: ECs on phase-scrambled images. We plot CIs for the EC to humans achieved by all models in the phase-scrambling corruption as an example, analogous to Figure 4
@@ 20 BiTM_resnetv2_101x3 46 alexnet
Table 2: Model Indices for MvH. We map model indices in Figure 4 to their full names.

### v2

@@ Appendix C Proof of Ceiling Effects
In Section˜2, we claim that as soon as one classifier achieves perfect accuracy on the investigated sequence of trials, the EC of the pair will be 0 irrespective of the accuracy of the other one. Here, we proof this claim, which is easiest to derive from a confusion matrix as shown in Table˜1. The observed agreement is found on the diagonal, $p_{obs}=a+d$, while the expected agreement is what would be on the diagonal if observers were independent ($p_{exp}=(a+b)(a+c)+(c+d)(b+d)$). For the proof, consider without loss of generality that classifier 2 was perfect, i.e. $a+c=1$.
@@
Appendix D Lower bound on EC
Figure 2b showed how the mismatch in accuracies between two classifiers imposes upper bounds on EC. In extension, Figure 6 shows that there are also non-trivial lower bounds (subplot b). These lower bounds are not a simple function of the upper bounds. Instead, they are inversely related such that when accuracies are inverted (i.e., $p_{1}=1-p_{2}$ along the off-diagonal) lower values can be attained. But perhaps counterintuitively, even with inverted accuracies, $\text{EC}=-1$ can not always be reached. This is in contrast to the upper bound, where aligned accuracies (i.e., $p_{1}=p_{2}$) always allows $\text{EC}=+1$.
Figure 6: Lower bound of EC in comparison to upper bounds. The left subplot (a) is the same as 2b, showing regions in which EC is upper bounded by the mismatch in accuracies two classifiers (x and y axes). Additionally, the right subplot (b) shows the lower bounds. Again, each point represents a model-vs-human comparison in the analysis of geirhos2021partial.
Appendix E Effect of EC on CI size
For Figure˜3, we have chosen a ground-truth EC of 0.5 to demonstrate how the EC between two classifiers depends on their accuracy and the number of trials on which EC is measured. However, the choice of the ground-truth EC influences the CI size as well, which we demonstrate in Figure˜7. We again generate data using our copy-model, and plot the ground-truth EC (input of the model) against the distribution of empirical ECs (calculated on model outputs). To better show the changes in CI size, we plot the delta between empirical and ground-truth EC rather than the absolute EC. Note how the CI bounds are not symmetric and largest at a ground-truth EC of 0.5.
@@ Appendix E Effect of EC on CI size
Appendix F Bootstrapping
For readers unfamiliar with bootstrapping, we provide a basic introduction here, but refer to efron1994introduction for details. The idea is the following: We want to estimate the uncertainty in a summary statistic (e.g., the mean), which is computed for a dataset $X$ of size $N$. If we had access to the underlying data-generating process, we could simply generate $M$ more datasets (each of size $N$) and compute the summary statistic for each. This would give us a distribution of values of the statistic, from which we could then compute a measure of uncertainty (e.g. a percentile interval of the distribution of means). Without access to the true data-generating process, we build the best possible parameter-free model: To generate a data point, randomly select one of the data points $x_{i}$ with uniform probability. To generate an entire dataset, draw $N$ data points with uniform probability, with replacement. In the context of error consistency between two observers, one data point corresponds to one trial, i.e. a pair of responses $r_{i}=(r^{(1)}_{i},r^{(2)}_{i})\in\{0,1\}^{2}$. The full algorithm to obtain a confidence interval for two sequences of responses $r^{(1)}$ and $r^{(2)}$ can thus be summarized in Algorithm˜1.
Algorithm 1 Bootstrapping EC confidence intervals
1:Arrays $r^{(1)},r^{(2)}\in\{0,1\}^{N}$, number of bootstrap iterations $M$, sample size NN
[…]
4:for $i=1$ to $M$ do
5:  $\mathcal{I}\leftarrow$ sample NN indices from $\{1,\ldots,N\}$ uniformly with replacement
6:  $ec\leftarrow\text{EC}(r^{(1)}[\mathcal{I}],r^{(2)}[\mathcal{I}])$ $\triangleright$ Compute error consistency
7:  $\mathcal{L}\leftarrow\mathcal{L}\cup\{ec\}$ $\triangleright$ Add to list
8:end for
[…]
10:return $\text{lower},\text{upper}]$
Similarly, the algorithm for estimating confidence intervals without access to real trials is described in Algorithm˜2.
Algorithm 2 Bootstrapping EC confidence intervals without real trials
1:Target marginals $P^{(1)},P^{(2)}$ over $K$ classes, target error consistency EC, number of trials NN, number of simulations $M$
[…]
3:Compute underlying parameters:
[…]
7:
[…]
9:for $i=1$ to $M$ do
10:  Generate responses for first classifier:
11:  for $j=1$ to NN do
12:   $r^{(1)}_{j}\sim\text{Categorical}(\tilde{P}^{(1)})$ $\triangleright$ Sample from marginal
13:  end for
14:
15:  Generate responses for second classifier:
16:  $N_{\text{copy}}\leftarrow\lfloor p_{\text{copy}}\cdot N\rfloor$ $\triangleright$ Number of responses to copy
17:  $\mathcal{I}_{\text{copy}}\leftarrow$ sample $N_{\text{copy}}$ indices from $\{1,\ldots,N\}$ without replacement
18:  for $j=1$ to NN do
19:   if $j\in\mathcal{I}_{\text{copy}}$ then
20:     $r^{(2)}_{j}\leftarrow r^{(1)}_{j} \triangleright$ Copy response
21:   else
22:     $r^{(2)}_{j}\sim\text{Categorical}(\tilde{P}^{(2)})$ $\triangleright$ Sample independently
23:   end if
24:  end for
25:
26:  $ec\leftarrow\text{EC}(r^{(1)},r^{(2)})$ $\triangleright$ Compute error consistency
27:  $\mathcal{L}\leftarrow\mathcal{L}\cup\{ec\}$
28:end for
[…]
30:return $\text{lower},\text{upper}]$
Figure 8: Convergence of EC estimates as a function of number of bootstraps. Starting from a single sequence of $1,000$ trials generated by our copy model (arbitrary target EC of $0.5$, classifier accuracies $0.7$ and $0.8$ respectively) we chose a number of bootstrap runs and bootstrapped $100$ times for each, to obtain estimates of uncertainty. Shaded regions correspond to 95% percentile intervals. Evidently, at the $M=10,000$ bootstrap runs we use, the estimates have stabilized, thus justifying our choice of $M$.
Appendix G Validating the copy model
As proven by Proposition˜2, our copy model is correct in the limit of trials. Additionally, we empirically validate the uncertainty estimates we derive from this model, by first generating some ground-truth data with a fixed error consistency at a fixed accuracy of both classifiers. We then bootstrap this data $M=1,000$ times, as we would to obtain percentile intervals. Next, we instead draw $M$ samples from our copy model, to compare the CIs implied by our copy model to those implied by the bootstrap, which we consider the gold standard. We plot this comparison in Figure˜9. Evidently, the distributions we obtain from our copy model match the bootstrapped results quite well, as long as the number of trials is sufficient or the accuracy avoids floor- and ceiling-effects.
Figure 9: Validating our copy model. We plot the distributions of EC estimates from the copy model to bootstrapped (gold standard) estimates of uncertainty. Top: 50 trials. Bottom: 1000 trials. Left: Accuracy 50%. Right: Accuracy 95%. Evidently, if both the number of trials is low and accuracy approaches the ceiling of 100%, the model diverges from bootstrapped results, but error consistency is not well-behaved in this case anyway. Ideally, experiments should avoid this regime, which can be achieved by increasing the number of trials and calibrating their difficulty.
Appendix H Details on Model-VS-Human results
@@
An idiosyncrasy of the model-vs-human benchmark that we have not expanded on so far is the fact that while the final scores (which are averages across the different corruptions) are stable, the uncertainty within each corruption can be quite large. In Figure˜11, we plot the CIs surrounding the EC values obtained for the phase-scrambling corruption as an example. In Figure˜12 we plot the pairwise rank correlations between all 17 corruptions, revealing that some, but not all corruptions are highly correlated, i.e. imply similar model rankings. In Figure˜10, we plot the confidence intervals one would obtain under the alternative aggregation strategy of simply averaging across conditions, without grouping by corruption first. Evidently, this would drastically affect the confidence intervals. To be clear, this is a deviation from the MvH protocol and thus “wrong”, but the data itself is the same, thus illustrating the point that there is large variance in the data, which is obfuscated by the hierarchical aggregation.
@@
In Figure˜2b, we explain how in many of the conditions within model-vs-human, the accuracy mismatch between models and human participants is so large that $k_{max}$ takes on very small values. One might wonder how model-vs-human scores would change if one excluded conditions not as prescribed by the authors, but based on this mismatch. However, this would render the selection of conditions dependent on the model. Hence, the benchmark would no longer present a fair comparison of models, and final scores would no longer be comparable.
Figure 10: EC to humans without hierarchical aggregation. If one were to simply define a model’s EC to humans as its average EC over all conditions (without first grouping by corruption), one would obtain the 95% confidence intervals depicted here, which clearly overlap for all models.
@@
Figure 11: ECs on phase-scrambled images. We plot CIs for the EC to humans achieved by all models in the phase-scrambling corruption as an example, analogous to Figure˜4
@@ 20 BiTM_resnetv2_101x3 46 alexnet
Table 2: Model Indices for MvH. We map model indices in Figure˜4 to their full names.
Appendix I Generating Synthetic Brain-Score Data
For the analysis conducted in Section˜4.2, we estimate the size of confidence intervals surrounding models on the public Brain-Score benchmark. To estimate a model’s EC to humans within one condition $c$, we would technically have to create one set of binary trials that has the property of having the desired average EC to the $n$ human observers of $\kappa_{real,c}$. To simplify, we drop this dependence and generate $n$ independent sequences of trials, one for each human who saw images in this condition. Generating only one sequence of trials would drastically overestimate the variance, which is reduced by taking the mean. For each of these sequences, we need three inputs for the copy-model: The desired kappa (which is given by the Brain-Score data after removing the ceiling correction), the accuracy of the humans (which is given by the model-vs-human data) and the accuracy of the model in this condition. The latter is unknown, but bounded mathematically. We obtain the CI sizes in Figure˜5 by selecting the accuracy in the middle of the bounds, which seems reasonable. Selecting the lower bounds of the accuracy leads to the most optimistic estimate of CI size, which we plot in Figure˜14.
Figure 14: Most optimistic Confidence Intervals for Brain-Score ECs. Like in Figure˜5, we plot the 95% confidence interval of the EC of the top 30 Brain-Score models, but assume the lowest mathematically possible model accuracy. The true CI size for these models probably looks more like Figure˜5, but this is a lower bound on their size.